Which has the greater area, a circle with a diameter of 4 feet or a square with a side of 4 feet?
Answer by edjones(8007)

Circle
A=pi*r^2
=pi*2^2
=12.566 ft^2
.
Square
A=s^2
=4^2
=16 ft^2
.
Ed